The Global Oscillation Network Group (GONG) Project is a community-based activity to operate a six-site helioseismic observing network, do the basic data reduction, provide the data and software tools to the community, and to coordinate analysis of the rich data set that is resulting. GONG data is available to any qualified investigator whose proposal has been accepted; however, active membership in a GONG Scientific Team encourages early access to the data and the collaborative scientific analysis that the Teams are undertaking. Information on the status of the Project, the scientific investigations, as well as access to the data is available on the WWW at www.gong.noao.edu.
Papers analyzing the tachocline at the base of the convection zone, helioseismic constraints on the solar radius, constraining the variation of the gravitational constant, and velocity flows around--and beneath--sunspots are among those recently submitted that will be highlighted on our URL.
The GONG network of instruments continues to demonstrate its reliability, obtaining more than 99% of all possible images during the second quarter of 1997. Of the images that were lost, about one-half resulted from two days of observations somehow vanishing without definitive cause. The data tape in question was in the primary bank and should have been collecting data from the time the tape was inserted until the secondary bank assumed writing data two days later. For reasons not yet understood, the data never made it onto the tape.
About 40% of the lost images were due to condensation forming on the turret optics during the winter months at Big Bear and El Teide. Big Bear suffered the most, and the problem has been addressed by replacing a faulty air dryer. The remaining 10% of the missing images were lost when a power supply breaker tripped at CTIO, and when we suffered failures of both Exabyte tape banks at El Teide. Thanks to the quick response and expertise of the on-site staff, the losses were kept to a minimum.
The reliability of the Exabyte tape drives has considerably improved since we discontinued using a defective batch of head-cleaning cartridges. The cleaning tapes used previously were contaminated with oil, and served only to degrade tape drive performance. Since we distributed the new cartridges, our headaches due to tape drive failures have decreased markedly.
The success of network operations has been due in large part to the efforts of the on-site staff at each of our sites. Their vigilance and willingness to "get their hands dirty" when needed is a major asset to the project.
During the past quarter, month-long (36-day) time series and power spectra were produced for GONG months 17, 18, and 19 (ending 970320) with fill factors of 0.92, 0.92, and 0.89. The fill factors for these three months were similar to those for the previous two months (0.91 and 0.94).
The preparation for the reprocessing campaign to regenerate p-mode power spectra from calibrated velocity images has been completed. During the quarter, the project temporarily halted the routine production of month-long p-mode power spectra and reprocessed GONG month 16 with the software and processing parameters that the project intended to use during the reprocessing campaign. After a thorough evaluation of the results and consultation with the project's scientific community through the DMAC Users' Committee, the project began reprocessing.
GONG months 17, 18, and 19 were processed for the first time and GONG months 15 and 16 have been reprocessed with the new p-mode processing parameters and software.
The Field Tape Reader (the subsystem that receives the raw data cartridges from the observing sites) processed 90 cartridges (104 during the previous quarter) containing 563 (606) site-days from the seven instruments. 378 (420) site-days were calibrated. The decrease in the number of site-days calibrated during the quarter is attributed to vacation schedules and the diagnosis and evaluation of a problem that was discovered in the calibration stage.
During the past quarter, the DSDS serviced 11 (13 during the previous quarter) data distribution requests for 3,286 (1,332) files totaling 0.9 (63.2) Gigabytes of data. Each of these data requests were filled during the day on which the request was received.The DSDS performed 1,532 (1,730) data cartridge transactions (library check-ins and check-outs) in response to requests from the data reduction pipeline and other internal operations.
The project has finished implementing the changes to the data processing as described in the previous NOAO Newsletter. The GONG DMAC Users Committee (DUC) met in Tucson on 18 June and approved the changes. Reprocessing is now proceeding both forwards and backwards in time from GONG Month 16. With this strategy the reprocessed time series will grow at the rate of two months per month.
Work continues in the area of spectral fitting, particularly in the areas of the asymmetrical line profile model, the leakage matrix calculation, the detailed modeling of the spectrum, and the development of alternative fitting methods. These are all substantial research efforts. Preliminary results from the wavelet de-noising and multi-taper package indicate that a spectrum produced with these methods will most likely result in improved estimates of the mode parameters.
Frederic Baudin and David Freilly-Fraillon have both now departed. Frederic has accepted a position at CfA, and David has returned to the Université de Nice. We wish them well! Jim Kennedy, GONG's project manager from its early years moved "upstairs" on 1 June to become the Operations Manager for the Gemini project. Jim will continue to be based in Tucson for at least the remainder of the calendar year. To effect a smooth transition, the Gemini project has generously agreed to let Jim continue to consult with GONG while he remains in Tucson. We'll all miss Jim. It is impossible to think of the GONG that we have achieved without his major contributions, and we wish him all the best in his new venture. It should be a piece of cake, it's "just" two sites right? Rob Hubbard has also left GONG. Rob performed a variety of critical roles for the project, beginning as an instrument design specialist, then moving to assistant project manager, and most recently operations manager. He leaves us to join the Breault Research Organization, a Tucson-based optics firm, so will still be "close."
While these changes punctuate GONG's transition from development, production, and deployment, to network operations, Jim and Rob leave in place a reliable, effective network of instruments, and a competent, enthusiastic staff poised to squeeze science out of the Sun for a solar cycle.
As operations are settling down, and the first major data reprocessing gets underway, we continue to make progress on "peak bagging" and on the development of a replacement camera system. As we get better and better frequency resolution data, it becomes clear that there are more and more things to take into account in the determination of the eigenfrequencies of the solar interior oscillations. The advancement of our techniques will clearly have to involve a renewed community participation. On the camera front, the good news is that a number of acceptable cameras have been identified. We also enjoyed an extremely productive "Red Team Review" of the camera replacement plan, with David Elmore, Sylvain Korzennik, Jesper Schou, Philip Stark, Steven Tomczyk, and Roger Ulrich asking a lot of hard questions, and helping to move that effort forward. The "proof of concept" effort should get started shortly as key positions get filled. Stay tuned!